# A Self-Supervised Learning Approach with Differentiable Optimization for UAV Trajectory Planning

### v1

@@ III-D4 Differentiable Optimization
Since the $\bm{\tau}^{*}$ is a function of $f(\bm{\theta}_{f})$ and $g(\bm{\theta}_{T})$, and will be passed to the upper level optimization for the final end-to-end training, the key part is to calculate the implicit gradient $\frac{\partial\bm{\tau}^{*}}{\partial\xi}$. The traditional unrolling approach maintain the computational graph throughout the entire iteration process, which poses significant computation burden when dealing with complex problem. It may also run into divergent or vanishing. In this work, we employ the implicit function differentiation theorem and leveraging the KKT condition of the Problem (4) at the optimal point to analytically compute the gradients of the parameters. Thus, there is no need for explicit unrolling of the entire iteration process. In practice, we leverage a fast differentiable QP solver for PyTorch as the optimization layer [38, 28].
@@
Figure 3: 3D ESDF cost map of the forest environment. (a) presents the point cloud reconstructed from collected depth images, while (b), (c), and (d) showcase the 3D ESDF map. We slice the ESDF map at different altitudes and color code the cost of the corresponding points.
@@
We develop a 3D cost map that provides traversability costs to guide UAV behavior by first reconstructing the environment offline from depth images, UAV poses, and camera poses. A 3D Euclidean Signed Distance Field (ESDF) like Figure 3 is then built based on the reconstructed environment. Unlike common approaches that assign distances only in obstacle regions (leaving free space as a constant), we also label the distance in free space to its nearest obstacle boundary. This ensures existence of valid gradient for network learning, as large internal volume of free space in the 3D space would otherwise cause gradient vanishing. Finally, we smooth the ESDF with a Gaussian filter to improve local differentiability. This 3D cost map enables self-supervision for our BLO framework.
@@ IV-1 Experimental Settings
For simulation, we use the open-source Autonomous Exploration Development Environment [10] and our customized UAV Gazebo simulator, running on a 2.4GHz i9 laptop with an NVIDIA RTX 4060 GPU. Real-world experiments are conducted on our custom-built quadrotor UAV with an NVIDIA Jetson Orin onboard computer running the planning pipeline and a Pixracer autopilot running our customized PX4 firmware. A front facing Intel RealSense D435 camera provides depth perception at 30 Hz. A motion capture system is used for indoor localization, which can be easily switched to visual inertial odometry [39, 40] for outdoor navigation.
@@ IV-2 Training Data
We collected training data from both simulated and real-world environments. In simulation, we used a joystick to manually fly the UAV in Gazebo simulator, collecting 3 datasets which contains approximately 2000 depth images and corresponding robot poses. We also collected 7 datasets in real-world to account for perception noise and varying lighting conditions. Notice that our method does not require labeled ground-truth data or human expert supervision; its training is fully self-supervised, relying solely on a well-designed loss and the 3D cost map.
@@ IV-3 Training Details
The perception network relies on a ResNet-18 as the backbone, and the planning network uses a combination of CNN and MLP with ReLU activation. The networks are trained with batch size of 16 using the Adam optimizer with a learning rate 0.0001. We used the pre-trained model from iPlanner and trained it for 50 epochs on an NVIDIA 4060 GPU. The entire training process took approximately 70 hours.
@@ IV-4 Baselines
We use the classic motion primitives planner (MP) as the state-of-the-art (SOTA) non-learning baseline. For the learning-based method, we adopt iplanner [25] as the baseline, which has already compared with other learning-based methods, including RL. Since iplanner only plans for 2D motion, we re-train it on our 3D cost map. All methods use a front-facing stereo vision depth camera.
@@ IV-A1 Overall Performance
We perform simulation under these environments: office, garage, and forest, see Figure 4. To evaluate the overall planning performance, we first measure the success rate, defined as the UAV reaching its goal from its starting point without any collision along the flight path. Specifically, we sample 60 start-goal pairs per environment. As shown in Table I, our method outperforms iPlanner in all cases because our differentiable MSTO module accounts for UAV dynamics in 3D, producing dynamically feasible trajectories. Although we notice the traditional MP performs slightly better in some scenarios, it can get stuck in local minima from which it cannot escape. For example, as shown in Figure 5, when the vehicle moves directly behind a pillar, its field of view is heavily restricted. In this scenario, target points located within a large area behind the obstacle often lead the MP planner to local minima. In contrast, our method demonstrates more robust, successfully generating collision-free and feasible trajectories.
@@ IV-A4 Inequality Constraints
Thanks to the iterative optimization in our MSTO, we can explicitly incorporate inequality constraints, such as flight corridor or actuator limit. Figure 6 shows a scenario where we incorporate corridor constraints. We can see our TO can successfully optimize the planned path within different required corridors. This capability is crucial for agile flight in constrained environments.
@@ IV-A5 Time Allocation
We finally evaluate our Time Allocation Net by comparing different time allocation strategies: uniform, acceleration–deceleration, and gradient descent with line search [37], see Table III. We can see that the uniform strategy is simple but ignores dynamic feasibility and environmental constraints, performing poorly in complex settings. The Acceleration-Deceleration strategy, usually built on a 5th-order polynomial, ensures smooth acceleration and deceleration with multiple initial and terminal conditions but may fail to optimize time allocation in obstacle-laden or dynamically constrained environments. Gradient Descent with line search cannot satisfy the real-time efficiency and is highly sensitive to the initial guess. In contrast, our approach incorporates iterative line search as the supervision to address these challenges, achieving optimality and computational efficiency for real-time execution.
@@ IV-B Real-World Experiment
To validate the real-time planning effectiveness of our method, we conducted real-world flight experiments in a complex environment using our custom-built quadrotor with an Intel RealSense D435 camera’s depth image as the only perception. The environment involves multiple vertical pillars, walls, horizontal beams and stacked boxes of varying height. Several narrow corridors are naturally generated between these obstacles. The depth measurements are also noisy during flight. Despite these challenges, our UAV successfully performed continuous obstacle avoidance maneuvers, demonstrating its ability to navigate dynamically in 3D space, as illustrated in Figure 7. The UAV exhibits smooth left-right and up-down evasive actions to avoid obstacles while maintaining stable trajectories. These results highlight the flexibility and robustness of our approach for obstacle avoidance in constrained 3D environments.
We also evaluate the overall tracking performance of our approach when avoiding one of the obstacles. As shown in Figure 8, it is clear that the tracking error of our approach is significantly lower than that of iPlanner during obstacle avoidance. Since iPlanner does not account for the dynamic feasibility of the planned path, it tends to have a larger tracking error when UAV conducting agile maneuver. Table IV presents a quantitative evaluation of these errors. Our approach achieves a mean tracking error of 0.0607m and a maximum error of 0.1268m. Moreover, the control effort is only 27.93m2/s7, all of which are significantly lower than those of iPlanner.
@@ V Conclusion
This paper develops a self-supervised UAV path planning pipeline that integrates a learning-based depth perception with differentiable trajectory optimization. A 3D cost map was introduced to self-supervise UAV behavior. Additionally, we designed a differentiable minimum snap trajectory optimization module to ensure dynamically feasible paths. A time allocation network improves the efficiency and optimality. Our approach thus improves generalizability and interpretability. Both simulation and real-world experiments demonstrate that our method can enable UAV navigate effectively by avoiding obstacles and execute dynamics feasible trajectory across various environments. Future work includes further testing under more complex environments and diverse operating conditions including dynamic obstacles and degraded lighting conditions. Our method achieves a 31.33% improvement in position tracking error and 49.37% reduction in control effort compared to the state-of-the-art.
Acknowledgement
We thank Xiangfu Li, Zemu Zhang and Xiangqin Chen for their assistance of hardware platform.

### v2

@@ III-D4 Differentiable Optimization
Since $\bm{\tau}^{*}$ is a function of $f(\bm{\theta}_{f})$ and $g(\bm{\theta}_{T})$, and will be passed to the upper level optimization for the final end-to-end training, the key part is to calculate the implicit gradient $\frac{\partial\bm{\tau}^{*}}{\partial\xi}$. The traditional unrolling approach maintain the computational graph throughout the entire iteration process, which poses significant computation burden. It may also run into divergent or gradient vanishing. In this work, we employ the implicit function differentiation theorem and leveraging the KKT condition of the Problem (4) at the optimal point to analytically compute the gradients of the parameters. Thus, there is no need for explicit unrolling of the entire iteration process. In practice, we leverage a fast differentiable QP solver for PyTorch as the optimization layer [1, 3].
@@
Figure 3: 3D ESDF cost map in the forest environment. (a) Reconstructed point cloud from depth images. (b–d) ESDF slices at different altitudes, color-coded by cost.
@@
We develop a 3D cost map that provides collision costs to guide UAV behavior by first reconstructing the environment offline from depth images, UAV poses, and camera poses. A 3D Euclidean Signed Distance Field (ESDF) like Figure 3 is then built based on the reconstructed environment. Unlike common approaches that assign distances only in obstacle regions (leaving free space as a constant), we also label the distance in free space to its nearest obstacle boundary. This ensures existence of valid gradient for network learning, as large internal volume of free space in the 3D space would otherwise cause gradient vanishing. Finally, we smooth the ESDF with a Gaussian filter to improve local differentiability. This 3D cost map enables self-supervision for our BLO framework.
@@ IV-1 Experimental Settings
For simulation, we use the open-source Autonomous Exploration Development Environment [6] and our customized UAV Gazebo simulator, running on a 2.4GHz i9 laptop with an NVIDIA RTX 4060 GPU. Real-world experiments are conducted on our custom-built quadrotor UAV with an NVIDIA Jetson Orin onboard computer running the planning pipeline and a Pixracer autopilot running our customized PX4 firmware. At runtime, trajectory optimizer outputs a time-parameterized sequence (position, velocity, acceleration) as the onboard outer-loop control reference. A front facing Intel RealSense D435 camera provides depth perception at 30 Hz. A motion capture system is used for indoor localization, which can be easily switched to visual inertial odometry [24, 25] for outdoor navigation.
@@ IV-2 Training Data
We collected training data from both simulated and real-world environments. In simulation, we used a joystick to manually fly the UAV in Gazebo simulator, collecting 3 datasets from the environments in Figure 4 (office, garage, and forest), which contains approximately 2000 depth images and corresponding robot poses. We also collected 7 datasets in real-world to account for perception noise and varying lighting conditions. Notice that our method does not require labeled ground-truth data or human expert supervision; its training is fully self-supervised, relying solely on a well-designed loss and the 3D cost map.
@@ IV-3 Training Details
The network follows an encoder–decoder architecture. A ResNet-18 backbone encodes visual observations, while the planning module fuses the encoded features with a linearly projected goal embedding. The fused representation is processed by two convolutional layers and a three-layer MLP to generate $k$ 3D waypoints, with a parallel MLP branch predicting a confidence score. The networks are trained with batch size of 16 using the Adam optimizer with a learning rate 0.0001. We used the pre-trained model from iPlanner [32] and trained it for 50 epochs on an NVIDIA 4060 GPU. The entire training process took approximately 70 hours.
@@ IV-4 Baselines
For the non-learning method, we choose two traditional approaches as baselines. MP [33]: Generate trajectories by searching over a set of pre-defined dynamically feasible motion segments. EGO-Planner [38]: A gradient-based planner without the ESDF construction. For the learning-based method, we adopt iPlanner [32] as the baseline, which has already compared with other learning-based methods, including RL. Since iPlanner only plans for 2D motion, we re-train it on our 3D cost map. All methods use a front-facing stereo vision depth camera.
@@ IV-A1 Overall Performance
We perform simulation experiments in Gazebo as shown in Figure 4. To evaluate the overall planning performance, we first measure the success rate, defined as the UAV reaching its goal from its starting point without any collision along the flight path. Specifically, we sample 60 start-goal pairs per environment. As shown in Table I, our method outperforms iPlanner in all cases because our differentiable MSTO module accounts for UAV dynamics in 3D, producing dynamically feasible trajectories. Although the traditional MP performs slightly better in some scenarios, it can get stuck in local minima from which it cannot escape. For example, as shown in Figure 5, when the vehicle moves directly behind a pillar, its field of view is heavily restricted. In this scenario, target points located within a large area behind the obstacle often lead the MP planner to local minima. In contrast, our method demonstrates more robust, successfully generating collision-free and feasible trajectories. We notice that EGO-Planner’s success rate is close to 100% under all environments, our method has more benefits on control effort (will discuss more in next section).
@@ IV-A4 Inequality Constraints
Thanks to the iterative optimization in our MSTO, we can explicitly incorporate inequality constraints, such as flight corridor or actuator limit. Figure 6 shows a scenario where we incorporate corridor constraints. Our method successfully optimizes the planned path within different required corridors. This capability is crucial for agile flight in constrained environments.
@@ IV-A5 Time Allocation
We finally evaluate our TAN by comparing different time allocation strategies: uniform, acceleration–deceleration, and gradient descent with line search [21], see Table III. We can see that the uniform strategy is simple but ignores dynamic feasibility and environmental constraints, performing poorly in complex settings. The Acceleration-Deceleration strategy, usually built on a 5th-order polynomial, ensures smooth acceleration and deceleration with multiple initial and terminal conditions but may fail to optimize time allocation in obstacle-laden or dynamically constrained environments. Gradient Descent with line search cannot satisfy the real-time efficiency and is highly sensitive to the initial guess. In contrast, our approach incorporates iterative line search as the supervision to address these challenges, achieving optimality and computational efficiency for real-time execution.
@@ IV-B Real-World Experiment
To validate the real-time effectiveness of our method, we conducted real-world flight experiments in a different physical room from where the training data was collected, with the depth camera as the only perception. The environment involves multiple vertical pillars, walls, horizontal beams and stacked boxes of varying height. Several narrow corridors are naturally generated between these obstacles. The depth measurements are also noisy during flight. Despite these challenges, our UAV successfully performed continuous obstacle avoidance, see Figure 7. The UAV exhibits smooth left-right and up-down evasive actions to avoid obstacles while maintaining stable trajectories. These results highlight the flexibility and robustness of our approach for obstacle avoidance in constrained 3D environments.
We also evaluate the overall performance of our approach when avoiding obstacles. Table IV shows the quantitative evaluation. Because all methods were deployed with the same low-level controller, tracking error directly reflects planning quality. Thanks to the MSTO, our approach achieves the lowest control effort with 27.93 m2/s7 while maintaining competitive tracking accuracy with the mean of 0.0607 m and maximum of 0.1268 m. Figure 8 illustrates a representative tracking run across all three methods.
@@ V Conclusion
This paper develops a self-supervised UAV path planning pipeline that integrates a learning-based depth perception with differentiable trajectory optimization. A 3D cost map was introduced to self-supervise UAV behavior. Additionally, we designed a differentiable minimum snap trajectory optimization module to ensure dynamically feasible paths. A time allocation network improves the real-time efficiency and optimality. Our approach thus improves generalizability and interpretability. Both simulation and real-world experiments demonstrate that our method can enable UAV navigate effectively by avoiding obstacles and execute dynamics feasible trajectory across various environments. Our method achieves 30.90% reduction in control effort compared to the state-of-the-art. Future work includes further testing under diverse operating conditions including dynamic obstacles and degraded lighting conditions.